1. A 9.37-L sample of gas has a pressure of 0.875 atm and a temperature of 81 °C. The sample is compressed to a volume of 8.20 L and is cooled to 20 °C. Calculate the new pressure of the gas, assuming that no gas escaped during the experiment. ____atm

2. A sample of CH4 gas has a volume of 367 mL, has a pressure of 704 mm Hg, and is at a temperature of 7.50 °C. Calculate the amount (in moles) of CH4 in the gas sample. ____mol

3. A sample of CCl4 with a pressure of 1.23 atm and a volume of 240 mL is allowed to react with excess CH4 at 127 °C.

CH4 (g) +  CCl4 (g) --> 2 CH2Cl2 (g)  

Calculate the pressure of the CH2Cl2 produced in the reaction if it is transferred to a 2.77-L flask and cooled to 28 °C.

_____atm
Answer #1

1)

Given:

Pi = 0.875 atm

Vi = 9.37 L

Vf = 8.20 L

Ti = 81.0 oC

= (81.0+273) K

= 354 K

Tf = 20.0 oC

= (20.0+273) K

= 293 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(0.875 atm*9.37 L)/(354.0 K) = (Pf*8.2 L)/(293.0 K)

Pf = 0.8276 atm

Answer: 0.828 atm

Only 1 question at a time please